Grade 10Trigonometry

What is the value of cos(22.5) in roots????? WITH FULL PROCESS

Ayush Kumar
9 Years agoGrade 10

1 Answer

Vaibhav Singh
9 Years ago
\cos45 = cos(2x22.5)
=> cos(2x22.5)  = \cos2(22.5) – \sin2(22.5)
= 2\cos2(22.5) -1 
=> 2\cos2(22.5) –1 = 1/\sqrt{2}
=>2\cos2(22.5) = (\sqrt{2} + 1)/\sqrt{2}
=> \cos2(22.5) = (\sqrt{2} + 1)/2\sqrt{2}
=> \cos(22.5) = \sqrt{(\sqrt{2} + 1)/2\sqrt{2}}